OpenStudy (anonymous):

i need help with this question factor fully 63x^2-48

OpenStudy (anonymous):

is this right 3(21x^2-16) ?

OpenStudy (anonymous):

nvrmind it is right

OpenStudy (badhi):

furthur more you can factorize as following \[\begin{align*}3(21x^2-16)=3(\sqrt{21}x-4)(\sqrt{21}x+4)\end{align*}\]

OpenStudy (anonymous):

but 21 isnt have a perfect square